# TPFL: Tsetlin-Personalized Federated Learning with Confidence-Based Clustering

### v3

@@ Abstract
The world of Machine Learning (ML) has witnessed rapid changes in terms of new models and ways to process users’ data. The majority of work that has been done is focused on Deep Learning (DL) based approaches. However, with the emergence of new algorithms such as the Tsetlin Machine (TM) algorithm, there is growing interest in exploring alternative approaches that may offer unique advantages in certain domains or applications. One of these domains is Federated Learning (FL), in which users’ privacy is of utmost importance. Due to its novelty, FL has seen a surge in the incorporation of personalization techniques to enhance model accuracy while maintaining user privacy under personalized conditions. In this work, we propose a novel approach called TPFL: Tsetlin-Personalized Federated Learning, in which models are grouped into clusters based on their confidence towards a specific class. In this way, clustering can benefit from two key advantages. Firstly, clients share only what they are confident about, resulting in the elimination of wrongful weight aggregation among clients whose data for a specific class may have not been enough during the training. This phenomenon is prevalent when the data are non-Independent and Identically Distributed (non-IID). Secondly, by sharing only weights towards a specific class, communication cost is substantially reduced, making TPLF efficient in terms of both accuracy and communication cost. The TPFL results were compared with 6 other baseline methods; namely FedAvg, FedProx, FLIS DC, FLIS HC, IFCA and FedTM. The results demonstrated that TPFL performance better than baseline methods with 98.94% accuracy on MNIST, 98.52% accuracy on FashionMNIST and 91.16% accuracy on FEMNIST dataset.
@@ 5.1 Phase A
6:Compute confidence score for each class $c$ on $D_{\text{conf}}$:
@@ 5.4 Phase D
Ultimately, the aggregator distributes the updated weight vectors $W_{\text{cluster}}^{k}$ to the corresponding clients, providing each client with the new global weights. The updated model is then evaluated on each client’s test data to assess performance improvements. Algorithm 2 shows Phase B, Phase C and Phase D in the aggregator.
@@
1:for each client $m$ do
2:     $k\leftarrow c_{max}$triangleright class $c_{max}$ determines cluster $k$ for $W_{m}$
3:     if weights $W_{m}^{{c}_{\text{max}}}$ are initial weights in cluster $k$ then
4:         Initialize cluster weights $W_{\text{cluster}}^{k}$ for cluster $k$:
@@
5:     else
6:         Aggregate received weights:
@@
7:     end if
8:end for
9:Compute average of weights in the cluster:
@@
10:Send updated weight vectors $W_{\text{cluster}}^{k}$ to corresponding clients of each cluster
11:Evaluate clients on $D_{\text{test}}$
5.5 High-Class Datasets
The communication costs among different clustering techniques are fundamentally different, specifically DL-based techniques compared with TPFL that relies on the TM algorithm. In case of high-class datasets, such as CIFAR100 or FEMNIST, TPFL has a key advantage in upload communication efficiency. This phenomenon stems from the fact that DL-based techniques necessitate the transmission of a full weight vector for the entire model, while in TPFL only the weight vector corresponding to the class with the highest confidence score is uploaded from each client device. In case of download communication cost, TPFL is still more efficient than other existing methods. However, assuming a federation of 100 clients whose data partitioning is based on CIFA100 dataset, the worst case scenario for the TPFL is that the download cost would be equal to the upload cost as no client in the federation shares a similar class to aggregate their weights. Therefore, no aggregation would take place and only weights are shared with the aggregator.
@@ 6.3 Data Partitioning
Both IID and non-IID data partition schemes are used in TPFL. Depending on different experimental setups, we used Dirichlet distribution to partition data into randomly shuffled IID and non-IID data. Dirichlet distribution uses a parameter $\alpha$ in order to define the severity of distribution. We used $\alpha=10000$ for IID data and $\alpha=0.05$ for non-IID data. Each TPFL client takes advantage of three different sets from to the end of training rounds; a training set for training the local model, a test set for the evaluation of local model’s performance and finally a confidence set so that the model’s confidence can be evaluated on.
@@ 6.6 Baselines
We chose the following methods in order to make a comparison between the existing methods and the proposed TPFL approach:
@@ 6.8 Comparison
Clients in the FL settings are normally evaluated under non-IID data partitioning schemes. Therefore, We compared the baseline methods with the proposed TPFL method under non-IID data partitioning. Hence, our comparison is between the best results of each baseline with non-IID data partitioning and TPFL under experiment 5. Table 6.7 shows the summary of our comparison. As it can be seen, the TPFL model outperforms all baselines in all datasets followed by IFCA and then FedTM. It is worth mentioning that TFPL implementation was carried out only in 10 rounds, the least number of rounds among all baseline methods. This phenomenon illustrates the positive effect of cluster-based personalization within TPFL setting. In addition, the communication cost of the TPFL in terms of uploaded and downloaded model parameters is the most optimized followed by FedTM, which itself is based on the TM algorithm.
This evident result shows the difference between DL-based models and the TM-based models. Based on the obtained results and their comparison with baseline models, it can be understood that the TPFL approach can be utilized in a one-shot learning fashion, in which each experimental setup is carried out only in one round. However, it is advisable to increase the number of local training epochs in order to make up for absence of multiple rounds of training.
@@ 7 Conclusion
With the advent of FL algorithms, numerous works have been put forward to maintain the users privacy. However, vanilla FL suffers from numerous disadvantages. PFL is considered a promising solution to solve the shortcomings of the vanilla FL. In this work, we propose a personalization method called Tsetlin-Personalized Federated Learning (TPFL) that leverages the voting mechanism of the TM in order to cluster clients based on their confidence towards a class. The results of TPFL compared with other FL and PFL baselines demonstrate that TPFL outperforms all baselines for MNIST, FashionMNIST and FEMNIST datasets. A promising future work that can be built upon TPFL is the utilization of innovative clustering strategies such as thresholding strategies for defining a confidence threshold and multiclass selection of class weights | sharing weights of more than one class | so that a client can be assigned to multiple clusters.

### v4

@@ Abstract
The world of Machine Learning (ML) has witnessed rapid changes in terms of new models and ways to process users’ data. The majority of work that has been done is focused on Deep Learning (DL) based approaches. However, with the emergence of new algorithms such as the Tsetlin Machine (TM) algorithm, there is growing interest in exploring alternative approaches that may offer unique advantages in certain domains or applications. One of these domains is Federated Learning (FL), in which users’ privacy is of utmost importance. Due to its novelty, FL has seen a surge in the incorporation of personalization techniques to enhance model accuracy while maintaining user privacy under personalized conditions. In this work, we propose a novel approach called TPFL: Tsetlin-Personalized Federated Learning, in which models are grouped into clusters based on their confidence towards a specific class. In this way, clustering can benefit from two key advantages. Firstly, clients share only what they are confident about, resulting in the elimination of wrongful weight aggregation among clients whose data for a specific class may not have been enough during the training. This phenomenon is prevalent when the data are non-Independent and Identically Distributed (non-IID). Secondly, by sharing only weights towards a specific class, communication cost is substantially reduced, making TPLF efficient in terms of both accuracy and communication cost. The TPFL results were compared with 5 other baseline methods; namely FedAvg, FedProx, FLIS, IFCA and FedTM. The results demonstrated that TPFL performance was better than baseline methods with 98.94% accuracy on MNIST, 98.52% accuracy on FashionMNIST and 91.16% accuracy on FEMNIST dataset.
@@ 5.1 Phase A
6:Compute and find confidence score for each class $c$ on $D_{\text{conf}}$:
@@ 5.4 Phase D
Ultimately, the aggregator distributes the updated weight vectors $W_{\text{cluster}}^{k}$ to the corresponding clients, providing each client with the new global weights. Each client is then evaluated on its test data to assess the performance improvements. Afterwards, the evaluation result of each model is received by the server. Algorithm 2 shows Phase B, Phase C and Phase D in the aggregator.
@@
1:for each round $r$ do
2:     for each client $m$ do
3:         $k\leftarrow c_{max}$triangleright class $c_{max}$ determines cluster $k$ for $W_{m}$
4:         if weights $W_{m}^{{c}_{\text{max}}}$ are initial weights in cluster $k$ then
5:              Initialize cluster weights $W_{\text{cluster}}^{k}$ for cluster $k$:
@@
6:         else
7:              Aggregate received weights:
@@
8:         end if
9:     end for
10:     for each cluster $k$ do
11:         Compute average of weights in the cluster:
@@
12:     end for
13:     Send updated weight vectors $W_{\text{cluster}}^{k}$ to corresponding clients of each cluster.
14:     Receive clients evaluation results on their $D_{\text{test}}$.
15:end for
5.5 TPFL Complexity
As it can be seen in algorithm 2, the entire federation runs for $R$ rounds, resulting in $O(R)$ time complexity. Each round consists of two main iterations. The first part is the training of $M$ clients, which has $O(RM)$ complexity. The second part is the calculation of the average weights of $K$ clusters, which has $O(RK)$ complexity. Therefore, the TPFL time complexity can be shown as $O(R(M+K))$.
@@ 6.3 Data Partitioning
Both IID and non-IID data partition schemes are used in TPFL. Depending on different experimental setups, we used Dirichlet distribution to partition data into randomly shuffled IID and non-IID data [33]. Dirichlet distribution uses a parameter $\alpha$ in order to define the severity of distribution. We used $\alpha=10000$ for IID data and $\alpha=0.05$ for non-IID data. Each TPFL client takes advantage of three different sets from to the end of training rounds; a training set for training the local model, a test set for the evaluation of local model’s performance and finally a confidence set so that the model’s confidence can be evaluated on.
@@ 6.6 Baselines
We chose 5 baseline methods in order to compare the proposed TPFL with. Table 3 demonstrates the specification summary of these baselines. These methods consist of both vanilla FL and PFL. Moreover, the table also shows non-IID data generation has increasingly favored the Dirichlet distribution in recent years. The implementation details of these methods is as follows:
@@
We chose a range of popular works that are based on both vanilla FL and PFL. In addition, we included both DL-based works and a TM-based method in order to show a general overview of the TM algorithm potential in FL settings. Given the TM numerous advantages, its comparison with DL-based methods in FL context can shed a light on TM future works.
Table 3: Summary of baseline methods specifications.
Baseline Year PFL Datasets Reimplementation Non-IID
FedAvg [34] 2016 No CIFAR-10, MNIST Full
LEAF
Library
FedProx [35] 2018 No MNIST Full
LEAF
Library
FLIS DC [19] 2022 Yes
CIFAR-100, CIFAR-10,
SVHN, FMNIST
Partial
Dirichlet
Distribution
FLIS HC [19] 2022 Yes
CIFAR-100, CIFAR-10,
SVHN, FMNIST
Partial
Dirichlet
Distribution
IFCA [12] 2020 Yes
MNIST, FEMNIST,
CIFAR-10
Partial NA
FedTM [29] 2023 No
MNIST, FMNIST,
FEMNIST
None
Dirichlet
Distribution
@@ 6.8 Comparison
Clients in the FL settings are normally evaluated under non-IID data partitioning schemes. Therefore, We compared the baseline methods with the proposed TPFL method under non-IID data partitioning. Hence, our comparison is between the best results of each baseline with non-IID data partitioning and TPFL under experiment 5. Table 6.7 shows the summary of our comparison. As it can be seen, the TPFL model outperforms all baselines in all datasets followed by IFCA and then FedTM. It is worth mentioning that TFPL implementation was carried out only in 10 rounds, the least number of rounds among all baseline methods. This phenomenon illustrates the positive effect of cluster-based personalization within TPFL setting.
This evident result shows the difference between DL-based models and the TM-based models. Based on the obtained results and their comparison with baseline models in figure 4, it can be understood that the TPFL approach can be utilized in a one-shot learning fashion, in which each experimental setup is carried out only in one round. However, it is advisable to increase the number of local training epochs in order to make up for absence of multiple rounds of training.
As it can be seen from table 6.7, in case of high-class datasets, such as FEMNIST, TPFL has a key advantage in upload communication efficiency per model. This phenomenon stems from the fact that DL-based techniques necessitate the transmission of a full weight vector for the entire model, while in TPFL only the weight vector corresponding to the class with the highest confidence score is uploaded from each local client. As shown in figure 5, in case of download communication cost, TPFL is still more efficient when it comes to communication cost. However, assuming a federation of 100 clients with CIFA100 dataset, the worst case scenario for the TPFL is that the download cost would be equal to the upload cost as no client in the federation shares a similar class to aggregate their weights. Figure 5 shows this correlation as the download communication cost progressively goes higher when clients have to deal with more classes of data.
@@ 7 Conclusion
With the advent of FL algorithms, numerous works have been put forward to maintain the users privacy. However, vanilla FL suffers from numerous disadvantages. Clustering-based PFL is considered a promising solution to address the shortcomings of the vanilla FL. Given the beneficial and explainable voting mechanism it has, the TM can be a great candidate for clustering-based PFL. In this work, we propose a personalization method called Tsetlin-Personalized Federated Learning (TPFL) that leverages the voting mechanism of the TM in order to cluster clients based on their confidence towards a class. The results of TPFL compared with other FL and PFL baselines demonstrate that TPFL outperforms all baselines for MNIST, FashionMNIST and FEMNIST datasets. A promising future work that can be built upon TPFL is the utilization of innovative clustering strategies such as thresholding strategies for defining a confidence threshold and multiclass selection of class weights | sharing weights of more than one class | so that a client can be assigned to multiple clusters.
8 Declarations
8.1 Competing Interests
The authors have no relevant financial or non-financial interests to disclose.
8.2 Authors Contribution Statement
Author Rasoul Jafari Gohari conceptualized and implemented the proposed method while all authors collaboratively structured the the article, prepared the manuscript, analyzed data, reviewed and edited drafts, and finally approved the publication.
8.3 Code availability
All the codes and implementations are available via our Github repository.
8.4 Funding
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
8.5 Compliance with Ethical Standards
All authors have followed the ethical standards in conducting and reporting this research, as outlined by the Springer Nature Code of Conduct for Authors.